Use the percent of sales method to forecast next years accounts payable. Current year sale are $24,500,000 and sales are expected to rise by 25%.
The firms accounts payable balance is $1,701,600. What is the projection for next years accounts payable?
A. 1,000,600
B. 2,127,000
C. 3,981,250
D. 6,125,000

Answer by gonzo(654)

this year's sales are:
24,500,000
next year's sales are 1.25 * this.
1.25 * 24,500,000 = 30,625,000
this year's accounts payables are:
1,701,600
this represents (1,701,600 / 24,500,000) * 100% of this year's sales:
(1,701,600 / 24,500,000) * 100% = .06945306122 * 100% = 6.945306122%
-----
6.945306122% of next year's sales = (6.945306122%/100%) * 30,625,000:
(6.945306122%/100%) * 30,625,000 = .06945306122 * 30,625,000 = 2,127,000
-----
answer is B
-----